# PKS 2131--021 – Discovery of Strong Coherent Sinusoidal Variations from Radio to Optical Frequencies: Compelling Evidence for a Blazar Supermassive Black Hole Binary

### v1

@@ 2 The Continuation of the 15 GHz Sine Wave Signal in PKS 2131--021
Note. — The period is that of the Haystack+OVRO data with the weights in the ratio of the number of data points ($879/154=5.71$) shown by the red sine wave in Fig. 2(b). This period is used as a prior in obtaining the parameters above for the best least squares sine wave fit to the OVRO data shown by the purple dashed curve in Fig. 1(a).
@@
Note. — $\mathcal{P}$ is the GLS power and $\Delta P$ is the range of periods included in the test. The tests are described in Sec. 2.1.
2.1 Statistical Significance of the Sinusoid Parameter Agreement in 1975–1981 vs. 2008–2023
@@
We carried out the generalized Lomb–Scargle (GLS) (Lomb, 1976; Scargle, 1982) periodogram analysis described in Section 3.4 of Paper 1. With this analysis we address the question whether a random process can spuriously produce the detected periodic signals in the OVRO and Haystack data. The Null hypothesis is that the light curves originate from a power-law red noise process that reproduces the power spectral density (PSD) and probability density function (PDF). We devise three tests.
@@
Test 1 – OVRO data: Under the Null hypothesis, what is the probability that the strongest peak in the GLS periodogram has a higher test statistic than the GLS peak of the the OVRO data? The test statistic is the significance of the GLS peak estimated at the peak frequency. We refer to Appendix A of Paper 1 for a detailed description of the procedure. Based on 300,000 simulations we estimate a probability of $=1.33\times 10^{-4}$ ($3.65\sigma$).
@@
Test 2 – Haystack data period: Under the Null hypothesis and knowing the period in the OVRO data, what is the probability that the strongest peak in the GLS periodogram of the Haystack data (a) has a power of at least 0.5 and (b) appears at a period in the $3\sigma$-uncertainty range spanning the period detected in the OVRO data? Based on 100,000 simulations we estimate a probability of $=1.97\times 10^{-2}$ ($2.06\sigma$).
@@
Test 3 – Haystack data phase: Under the Null hypothesis, and assuming the Haystack data shows a periodicity with period matching that of the OVRO data, what is the probability that the phases match to within 20% of a half-period? In the case of a randomly produced spurious periodicity, the phase would be uniformly distributed. Therefore, the probability for test 3 is 0.2.
@@
We estimate the probability of all three observations as defined in tests 1–3 under the Null hypothesis by multiplying the probabilities of the three independent tests. The joint probability is $4.25\times 10^{-7}$ ($4.88\sigma$). All tests and results are summarized in Table 2.
@@
We reject the Null hypothesis at $>3\sigma$ significance level based on the OVRO data alone, and with much higher confidence based on the joint results from the OVRO and Haystack data analysis. It is highly unlikely that a random process spuriously produced the observed periodicity. Consequentially, a different, non-random process must have produced the periodic signal.
@@ 3 The Radio Spectrum, Structure, and Polarization of PKS 2131--021
The polarization image of Fig. 4 is produced by first averaging the total intensity $I$ and Stokes parameters $Q$ and $U$ over the 16 observing epochs between 1997 and 2012, and then calculating the fractional polarization and electric vector position angle (EVPA) images from the stacked quantities (see Pushkarev et al., 2023, for details). Stacking the polarization this way reveals the underlying magnetic field structure. It also reveals a much broader cross-section of the entire jet, which is not often fully visible at individual epochs.
@@
Figure 5: Polarization of PKS 2131--021 observed in ALMA Band 3 (Cyan points: 91.5 GHz; green points: 103.5 GHz.). The EVPA changes little, if at all, between 15 GHz and 345 GHz, and shows no dramatic shift between the optically thick and optically thin regimes.
The integrated linear polarization in PKS 2131--021 stays predominantly parallel to the jet direction all the way up to 350 GHz (Rocha et al., 2023). As described in section 4.1, PKS 2131-021 is an ALMA calibrator, and we have used that data extensively in this paper. The variation in the linear polarization at 91.5 GHz and 103.5 GHz from 2019 to 2023.5 seen by ALMA is shown in Fig. 5, where it is seen that the EVPA changes little throughout the sinusoidal cycle.
@@
Figure 6: PKS 2131--021 light curves at radio, mm, sub-mm, and optical wavelengths showing a steady monotonic progression of the phase of the sinusoidal variations with frequency. (a) Haystack 2.7–31.4 GHz observations (O’Dea et al., 1986) together with the least-squares sine-wave fits to the data at each of the frequencies. There is a steady progression in phase with frequency between the four light curves, with the low-frequency sine waves lagging the high-frequency sine waves. (b) OVRO 15 GHz light curve plus ALMA light curves at 91.5, 104, and 345 GHz. As in (a), there is a continuing progression in the phases of the sine wave fits with frequency. (c) Comparison of the ZTF optical and OVRO radio light curves of PKS 2131--021. The ZTF $i$, $r$, and $g$-band light curves of PKS 2131--021 are shown. The large symbols are average values for each year. The black points show the OVRO 15 GHz points over this period, which provide the phase reference point. The curves show the least squares sine wave fits to the corresponding data. We associate the optical peak with the nearest radio peak, not the preceding radio peak. (d) The observed phase shifts, relative to the OVRO 15 GHz light curve, of the sine-wave fits to the data, given in Table 3. A positive phase shift indicates that the light curve is shifted to a later time than the 15 GHz light curve. The curve shows a quadratic polynomial fit to the phase offset vs. log(frequency) data (see Sect. 4.3).
Table 3: Phase Shift with Frequency Relative to the 15 GHz Light Curve
Frequency Phase Uncertainty
Band (Fraction (Fraction
of a cycle) of a cycle)
2.7 GHz 0.147 0.03
7.9 GHz 0.048 0.015
15 and 15.5 GHz 0 0
31.4 GHz $-0.035$ 0.015
91.5 GHz $-0.072$ 0.01
104 GHz $-0.073$ 0.01
345 GHz $-0.135$ 0.01
infrared $-0.27$ 0.05
Optical $-0.37$ 0.01
@@ 4 The Centimeter-to-Optical Emission in PKS 2131--021
4.1 The Centimeter-to-Submillimeter Emission
The Haystack observations from 1975 to 1983 at 2.7, 7.9, 15.5, and 31.4 GHz (O’Dea et al., 1986) are shown in Fig. 6(a), and the OVRO and ALMA lightcurves between 2008 and 2023 at 15, 91.5, 104, and 345 GHz are shown in Fig. 6(b). The ALMA data were downloaded from the ALMA calibrator website. The color-coded curves are least-squares fits of sine waves to the data, derived by the procedure presented in Paper 1. The phase offsets derived from the sine-wave fits of Figs. 6(a) and (b) are given in Table 3. The 15.5 GHz (Haystack) and 15 GHz (OVRO) light curves were used as reference light curves, since these frequencies occur in the earlier lower-frequency and the later higher-frequency data sets, respectively, and these frequencies are essentially the same to the accuracy needed here. Note that the phase shifts vary monotonically with frequency from the lowest to the highest frequency of observation, i.e. from 2.7 GHz to the optical band. To the best of our knowledge, such coherence in the flux density variations in a blazar, extending from the radio band to the optical band, has not been observed before.
4.2 The Infrared to Optical Emission in PKS 2131--021
@@
4.2.2 Strong Optical Sinusoidal Emission
@@
The ZTF $g$, $r$, and $i$ band light curves of PKS 2131--021 (Fig. 6c) show a periodic variation having the same period, to within the uncertainties, as the 15 GHz light curve. Sine-wave fits to the ZTF data are presented in Appendix C. where we show that the combined sine-wave fit to all three data sets yields a period of $P_{\textrm{optical}}=(1764\pm 36)$ days. This agrees, within the uncertainties, with the period of $(1739.2\pm 1.2)$ days determined from the OVRO+Haystack data. It appears, therefore, that the sinusoidal variations persist at least up to optical wavelengths. We present evidence below that the same periodicity likely continues up to $\gamma$-ray energies. The time shift between the observed maxima in the sine wave fits to the optical light curves and the 15 GHz curve is 638 or 1101 days, depending on the direction of the shift. The more obvious choice is 638 days, since this shifts the light curve in the same direction relative to the 15 GHz OVRO light curve as we determined from cm to sub-mm wavelengths, and continues the monotonic trend we have observed at the lower frequencies (see Fig. 6(d)). We return to this point in section 6.
@@
4.2.3 The Comparative Difficulty of Detecting Optical vs. Radio Sinusoidal Variations in Blazars
This is the first detection of statistically compelling sinusoidal variations in a blazar at both optical and radio wavelengths. It provides the opportunity to estimate the relative difficulty of making such detections in these two windows. As shown in Paper 1, the peak random fractional variations in the radio light curve are $\sim 20\%$ (see Fig. 10(c) of Paper 1). From the ZTF data of Fig. 6(c) we find that the peak random fractional variations in the optical light curve are $\sim 200\%$. This order of magnitude difference, coupled with the extraordinary challenges of determining the high statistical significance of the sinusoidal variations in PKS 2131--021, with all the various types of confusion that are introduced by the random variations discussed in Paper 1 and the inevitable annual gaps in the optical light curves, reveals that the detection of statistically significant optical sinusoidal variations in blazars is a daunting challenge. The situation is illustrated in Fig. 7. There are periods when the optical short-term variations are no larger than those at radio wavelengths, but there are also periods when they are an order of magnitude, or more, greater than those at radio wavelengths, making it more difficult to detect sinusoidal variations at optical than at radio wavelengths.
@@
In order to obtain a qualitative measure of the relative difficulty of detecting a blazar showing sinusoidal variations like those seen in PKS 2131--021 at optical and radio wavelengths, we have carried out the same GLS analysis on the optical data as on the radio data, using the optical data shown in Fig. 7. We found that the local $p$-values of the GLS peaks in the $i$, $r$, and $g$ bands are 0.5, 0.4, and 0.5, respectively, whereas that at 15 GHz is 0.002.
@@ 4.3 Coherence of the Periodicity from the Radio to the Optical Emission
$\delta\phi=0.024-0.125\log_{10}(\nu/15)$ (1)
where $\nu$ is in GHz. We make use of this relationship in equation (3) below. The curved line of Fig. 6(d) shows a quadratic polynomial fit to all the data given by $\delta\phi=0.176-0.144\;\log_{10}(\nu/15)+0.0087[\log_{10}(\nu/15)]^{2}$.
@@ 5 X-ray Observations and Future Prospects
Hard X-ray emission from blazars is important because it is directly connected to the central engine. The soft X-ray emission from blazars is thought to be due to synchrotron emission, while the hard X-ray emission is thought to be caused by up-scattering of soft photons either from the synchrotron emission (Mastichiadis & Kirk, 2002), or from relatively cool structures like the accretion disk and the broad line region (Dermer & Schlickeiser, 1993; Sikora, 1994). While significant periodicities are common in the light curves of stellar mass black hole X-ray binary systems, they are surprisingly rare in AGN X-ray observations (however, see Juryšek et al. 2022). Only very recently have significant periodic signals started to be widely reported and studied systematically (Arcodia et al., 2021; Miniutti et al., 2019). In most cases, they originate in stellar-mass compact objects interacting with the AGN accretion disk (Franchini, Alessia et al., 2023). In the particular case of blazars, the X-ray light curves are usually chaotic and aperiodic. X-ray spectral-timing studies have revealed the existence of time-dependent flux density and spectral state patterns. Bhatta et al. (2018) explored the connections between blazar X-ray variability and other properties, such as spectral hardness and intrinsic X-ray flux in the NuSTAR band. All 13 sources in their sample displayed high-amplitude, rapid, aperiodic variability with a timescale of a few hours.
Since X-rays probe the region close to the central engine, the presence of a close-separation SMBHB could potentially imprint the high-energy emission. Simulations predict that close-separation accreting SMBHBs will have different X-ray spectra to those of single accreting SMBHs (Tang et al., 2018; Krolik et al., 2019). Accreting SMBHBs may show a periodically modulated hard X-ray component whose period is of order the binary orbital period, although these predictions have been challenged (e.g., see Saade et al., 2023, 2024). There are no available high-quality X-ray data for PKS 2131--021, but short exposures during an XMM-Newton slew indicate a fairly high X-ray flux ($\sim 2\times 10^{-12}\rm~{}erg\,s^{-1}\,cm^{-2}$) in the 2–12 keV range. We analyzed all available Swift/XRT obervations. The stacked spectrum shown in Fig. 9 reaches a total net exposure time of 20.7 ks, characterized by a featureless power-law ($\Gamma=1.78\pm 0.15$) with a moderate intrinsic hydrogen column density of $N_{\mathrm{H}}=3.4^{+2.7}_{-2.4}\times 10^{21}\,\rm cm^{-2}$ at the rest-frame of $\rm z=1.283$. We estimate the 0.3–10 keV unabsorbed flux in the observed frame as $(1.86\pm 0.15)\times 10^{-12}\,\rm erg\,s^{-1}\,cm^{-2}$. In the near future, long-term monitoring with large-field-of-view instruments such as eRosita and Swift will be practical and promising, given the relatively high redshift of PKS 2131--021.

### v2

@@ 1. Fit the lightcurve at each frequency as a scaled and offset version of the OVRO light curve, or
Note. — The fitted parameters are explained in Sec. 2.1. The period is that of the weighted Haystack+OVRO data (see text) shown by the red sine wave in Fig. 2(b). This period is used as a prior in obtaining the parameters above for the least squares sine wave fit to the OVRO data shown by the purple dashed curve in Fig. 1(a).
Table 3: Probabilities and Significance Levels of GLS Tests from Simulations with Matched Red Noise Tail
Test Test GLS $\mathcal{P}_{\rm peak}$ Period Range ($\Delta P$) Total Number that $p$-value Significance
Number (max=1) (days) Simulations pass test ($\sigma$)
1 OVRO 0.83 All $10^{6}$ 53 $5.3\times 10^{-5}$ 3.88
2 Haystack $\geq 0.50$ 1657.5 – 1854.1 10000 197 $1.97\times 10^{-2}$ 2.06
3 phase - - - - 0.2 $0.84$
- 1+2+3 - - - - $2.09\times 10^{-7}$ $5.06$
@@
Note. — $\mathcal{P}$ is the GLS power and $\Delta P$ is the range of periods included in the test. The tests are described in Sec. 4.
4 Statistical Significance of the Sinusoid Parameter Agreement in 1975–1981 vs. 2008–2023
@@
We carried out the generalized Lomb–Scargle (GLS) (Lomb, 1976; Scargle, 1982; Zechmeister & Kürster, 2009) periodogram analysis described in Section 3.4 of Paper 1. With this analysis we address the question of whether a random process can produce the detected periodic signals in the OVRO and Haystack data. The null hypothesis is that the light curves originate from a power-law red noise process that reproduces the power spectral density (PSD) and probability density function (PDF). We model the PSD with a frequency ($\nu$) dependent power-law $\propto\nu^{-\beta}$ with index $\beta=2.05$. We model the PDF through the empirical cumulative distribution function of the original light curve. We refer to Appendix A in Paper 1 for details. We devised three tests, the results of which are given in Table 3:
@@
Test 1 – OVRO data: Under the null hypothesis, what is the probability that the strongest peak in the GLS periodogram has a higher test statistic than the GLS peak of the the OVRO data? The test statistic is the significance of the GLS peak estimated at the peak frequency. We refer to Appendix A of Paper 1 for a detailed description of the procedure. Based on $10^{6}$ simulations we estimate a probability of $5.3\times 10^{-5}$ ($3.88\sigma$).
@@
Test 2 – Haystack data period: Under the null hypothesis and knowing the period in the OVRO data, what is the probability that the strongest peak in the GLS periodogram of the Haystack data (a) has a power of at least 0.5 and (b) appears at a period in the $3\sigma$-uncertainty range spanning the period detected in the OVRO data? Based on 10,000 simulations we estimate a probability of $1.97\times 10^{-2}$ ($2.06\sigma$).
@@
Test 3 – Haystack data phase: Under the null hypothesis, and assuming the Haystack data shows a periodicity with period matching that of the OVRO data, what is the probability that the phases match to within the observed 20% of a half-period offset seen between the Haystack and OVRO sine wave fits? In the case of a randomly produced periodicity, the phase would be uniformly distributed. Therefore, the probability for test 3 is 0.2 (Paper 1).
@@
We estimate the probability of all three observations as defined in tests 1–3 under the null hypothesis by multiplying the probabilities of the three independent tests. The joint probability is $2.09\times 10^{-7}$ ($5.06\sigma$).
@@
We reject the null hypothesis at $>3\sigma$ significance level based on the OVRO data alone, and with much higher confidence based on the joint results from the OVRO and Haystack data analysis. It is highly unlikely that a random process produced the observed periodicity. Consequentially, a non-random process must have produced the periodic signal.
@@ 5 The Radio Spectrum, Structure, and Polarization of PKS 2131--021
The polarization image of Fig. 5 is produced by first averaging the total intensity $I$ and Stokes parameters $Q$ and $U$ over the 16 observing epochs between 1997 and 2012, and then calculating the fractional polarization and electric vector position angle (EVPA) images from the stacked quantities (see Pushkarev et al., 2023, for details). Stacking the polarization this way reveals the underlying magnetic field structure. It also reveals a much broader cross-section of the entire jet, which is not often fully visible in individual epochs.
@@
Figure 6: Polarization of PKS 2131--021 observed in ALMA Band 3 (Cyan points: 91.5 GHz; gray points: 103.5 GHz.). The EVPA changes little, if at all, between 15 GHz and 345 GHz, and shows no dramatic shift between the optically thick and optically thin regimes.
The integrated linear polarization in PKS 2131--021 stays predominantly parallel to the jet direction all the way up to 350 GHz (Rocha et al., 2023). The variation in the linear polarization at 91.5 GHz and 103.5 GHz from 2019 to 2023.5 seen by ALMA is shown in Fig. 6, where it is seen that the EVPA changes little throughout the sinusoidal cycle.
@@ 6 The Centimeter-to-Optical Emission in PKS 2131--021
In Fig. 3 (a), (b), and (d), we saw that the phase shift varies monotonically with frequency from 2.7 GHz to 345 GHz. In this section we show that this coherent phase shift with frequency continues through the infrared to the optical band. To the best of our knowledge, such coherence in the flux density variations in a blazar, extending from the radio band to the optical band, has not been observed before. We suggest that this phase shift with frequency is due to optical depth effects in the jet.
@@
6.2 The Strong Optical Sinusoidal Emission
@@
The ZTF $g$, $r$, and $i$ band light curves of PKS 2131--021 (Fig. 3c) show a periodic variation having the same period, to within the uncertainties, as the 15 GHz light curve. Sine-wave fits to the ZTF data are presented in Appendix A. where we show that the combined sine-wave fit to all three data sets yields a period of $P_{\textrm{optical}}=(1764\pm 36)$ days. This agrees, within the uncertainties, with the period of $(1739.2\pm 1.2)$ days determined from the OVRO+Haystack data. It appears, therefore, that the sinusoidal variations persist at least up to optical wavelengths. We present evidence below that the same periodicity likely continues up to $\gamma$-ray energies.
To determine the phase shift of the optical light curves relative to the OVRO 15 GHz light curve we have used the period $P=1739.2$ days derived for the Haystack 15.5 GHz and OVRO 15 GHz light curves in §2.1. The reasons for this choice are (i) the period of the optical sine wave is well determined, as can be seen from Fig. 3(c), (ii) it is consistent with the period derived from the OVRO+Haystack light curves, and furthermore (iii) in Fig. 7, where we compare the OVRO 15 GHz light curve with the ZTF light curves, it is immediately clear that the non-sinusoidal variations in the optical light curves are not seen at 15 GHz. The derived value of the optical sinusoidal shift is given in Table 2.
The time shift between the observed maxima in the sine wave fits to the optical light curves and the 15 GHz curve is 612 or 1127 days, depending on the direction of the shift. The more obvious choice is 612 days, since this shifts the light curve in the same direction relative to the 15 GHz OVRO light curve as we determined from cm to sub-mm wavelengths, and continues the monotonic trend we have observed at the lower frequencies (see Fig. 3(d)). We return to this point in section 8.
@@
6.3 The Comparative Difficulty of Detecting Optical vs. Radio Sinusoidal Variations in Blazars
This is the first detection of statistically compelling sinusoidal variations in a blazar at both optical and radio wavelengths. It provides the opportunity to estimate the relative difficulty of making such detections in these two windows. As shown in Paper 1, the peak random fractional variations in the radio light curve are $\sim 20\%$ (see Fig. 10(c) of Paper 1). From the ZTF data of Figs. 3(c) we find that the peak random fractional variations in the optical light curve are $\sim 200\%$. This order of magnitude difference, coupled with the extraordinary challenges of determining the high statistical significance of the sinusoidal variations in PKS 2131--021, with all the various types of confusion that are introduced by the random variations discussed in Paper 1 and the inevitable annual gaps in the optical light curves, reveals that the detection of statistically significant optical sinusoidal variations in blazars is a daunting challenge. The situation is illustrated in Fig. 7. There are periods when the optical short-term variations are no larger than those at radio wavelengths, but there are also periods when they are an order of magnitude, or more, greater than those at radio wavelengths, making it more difficult to detect sinusoidal variations at optical than at radio wavelengths.
@@
In order to obtain a qualitative measure of the relative difficulty of detecting a blazar showing sinusoidal variations like those seen in PKS 2131--021 at optical and radio wavelengths, we have carried out the same GLS analysis on the optical data as on the radio data, truncated to the ZTF time frame, using the optical data shown in Fig. 7. Based on 20,000 simulations, we found that the global $p$-values of the GLS peaks in the $i$, $r$, and $g$ bands are 0.5, 0.4, and 0.5, respectively, whereas that at 15 GHz is 0.002.
@@ 6.4 Coherence of the Periodicity from the Radio to the Optical Emission
$\delta\phi=0.020-0.116\log_{10}(\nu/15)$ (2)
where $\nu$ is in GHz. We make use of this relationship in equation (4) below. The curved line of Fig. 3(d) shows a quadratic polynomial fit to all the data given by $\delta\phi=0.178-0.146\;\log_{10}(\nu)+0.0093[\log_{10}(\nu)]^{2}$.
@@ 7 X-ray Observations and Future Prospects
Hard X-ray emission from blazars is important because it is directly connected to the central engine (scale being smaller than gravitational radius, Bhatta et al., 2018). The soft X-ray emission from blazars is thought to be due to synchrotron emission, while the hard X-ray emission is thought to be caused by up-scattering of soft photons either from the synchrotron emission (Mastichiadis & Kirk, 2002), or from relatively cool structures like the accretion disk and the broad line region (Dermer & Schlickeiser, 1993; Sikora, 1994). While significant periodicities are common in the light curves of stellar mass black hole X-ray binary systems, they are surprisingly rare in AGN X-ray observations (however, see Juryšek et al. 2022). Only very recently have significant periodic signals started to be widely reported and studied systematically (Arcodia et al., 2021; Miniutti et al., 2019). In most cases, they originate in stellar-mass compact objects interacting with the AGN accretion disk (Franchini, Alessia et al., 2023). In the particular case of blazars, the X-ray light curves are usually chaotic and aperiodic. X-ray spectral-timing studies have revealed the existence of time-dependent flux density and spectral state patterns. Bhatta et al. (2018) explored the connections between blazar X-ray variability and other properties, such as spectral hardness and intrinsic X-ray flux in the NuSTAR band. All 13 sources in their sample displayed high-amplitude, rapid, aperiodic variability with a timescale of a few hours.
Since X-rays probe the region close to the central engine, the presence of a close-separation SMBHB could potentially imprint the high-energy emission. Simulations predict that close-separation accreting SMBHBs will have different X-ray spectra to those of single accreting SMBHs (Tang et al., 2018; Krolik et al., 2019). Accreting SMBHBs may show a periodically modulated hard X-ray component whose period is of order the binary orbital period, although these predictions have been challenged (e.g., see Saade et al., 2023, 2024). There are no available high-quality X-ray data for PKS 2131--021, but short exposures during an XMM-Newton slew indicate a fairly high X-ray flux ($\sim 2\times 10^{-12}\rm~{}erg\,s^{-1}\,cm^{-2}$) in the 2–12 keV range. We analyzed all available Swift/XRT obervations. The stacked spectrum shown in Fig. 9 reaches a total net exposure time of 20.7 ks, characterized by a featureless power-law ($\Gamma=1.78\pm 0.15$) with a moderate intrinsic hydrogen column density of $N_{\mathrm{H}}=3.4^{+2.7}_{-2.4}\times 10^{21}\,\rm cm^{-2}$ at the rest-frame of $z=1.283$. We estimate the 0.3–10 keV unabsorbed flux in the observed frame as $(1.86\pm 0.15)\times 10^{-12}\,\rm erg\,s^{-1}\,cm^{-2}$. In the near future, long-term monitoring with large-field-of-view instruments such as eRosita and Swift will be practical and promising, given the relatively high redshift of PKS 2131--021.